Grade 12th passMechanics

A balloon of weight W descends with a acceleration f m/s2. the weight that must be thrown out in order give the baloon an equal upward acceleration will be?

mohan
11 Years agoGrade 12th pass

1 Answer

Vivek Kumar
9 Years ago
Let the upthrust due to air be P. Now, W - P = (W/g)fP - (W - g) = (W - Q)f/gAdding, Q/f = W/g + (W - Q)/g=> Q = 2Wf/(g + f)